OpenStudy (anonymous):

Describe how you can tell whether two lines are parallel, perpendicular, or neither without graphing them.

OpenStudy (ddcamp):

Look at the slopes (m) If the slopes are equal, they are parallel. If the slopes are "negative reciprocals," they are perpendicular. Negative reciprocal: \[m \rightarrow -\frac{ 1 }{ m }\]

OpenStudy (anonymous):

or another way to say what DDCamp said is, let m_1 be the slope of the first line, and let m_2 be the slope of the second line. if \[m_1*m_2=-1\], then they are parallel

OpenStudy (ddcamp):

@DemolisionWolf That would be perpendicular, just a typo I'm assuming.

OpenStudy (anonymous):

haha my bad

OpenStudy (anonymous):

So how do you know if it's neither?

OpenStudy (anonymous):

", then they are perpendicular"

OpenStudy (anonymous):

if m_1 = m_2, they are parallel if m_1*m_2=-1, they are perpendicular if neither is the case, then they are just two lines that will intersect somewhere.

OpenStudy (zpupster):

or the same exact line like 12=2x+4 6=x+2

OpenStudy (anonymous):

Well thank you @DemolisionWolf @DDCamp @zpupster c:

OpenStudy (zpupster):

parallel no solutions perpendicular 1 solution same line infinite no of solutions

OpenStudy (anonymous):

zpupster is right too, tons of good info on this post for you ^_^

OpenStudy (anonymous):

Oh goodness, thank you, this has helped a lot cx

OpenStudy (ddcamp):

Though with @zpupster's method, you have to be careful. Lines that aren't perpendicular also have 1 solution to the system (if they're not parallel or the same).

OpenStudy (anonymous):

Oh well, I look out for that c:

OpenStudy (anonymous):

I'll*

OpenStudy (zpupster):

true